the nuclear spin quantum number for deuterium, 2H is I=1. The number of nuclear spin states is ____?
Answer #1

Ans.

The number of nuclear spin states of a nucleus is 2I + 1 where I is the nuclear spin quantum number

For Deuterium , 2H , I = 1

So , number of nuclear spin states of Deuterium , 2H = 2I + 1 = 2*1 + 1 = 2 + 1 = 3

Hence , number of nuclear spin states of Deuterium , 2H is 3